Confusinated:

Graphing Linear Equations in 2 Variables

Confusinated:

Confusinated:

@JustSaiyan

Confusinated:

@Elsa213

Confusinated:

@Jason64

Angle:

Let's try plugging in x = 4 into the given equation what would you get for \(y\)?

Angle:

oh my bad, I mean to say plug in x = 2 that would be easier

Confusinated:

actually angle i have no idea how to do this at all so yea my teacher just gave it to me

Angle:

ok, no problem here's an easy example of what I mean by "plug in" say we have the equation \(y = 2x\) this is basically "I double whatever you give me" so if you give me \(2\) (for x), I give you \(4\) (as y) essentially, when you give me "2", plugging in looks like \(y = 2(2) = 4\)

Confusinated:

oh it make a little more sense but i am still confussed

Angle:

So when we plug in \(x = 2\) into the given equation \(-4y = x -18\) we stick the \(2\) in like this: \(-4y = (2) - 18\) and we can "solve" for y like this \(-4y = -16\) (then divide both sides by negative 4) and get \(y = 4\)

Confusinated:

aww okay

Angle:

so one of the points you can graph is: when x=2 you have y=4 you may also see this in class as (x, y) = (2, 4) do you know how to graph coordinates? :)

Confusinated:

kind of

Confusinated:

well i got the answer thank you very much

Angle:

no problem, I'm glad I could help if you need any more explaining, feel free to ask :)

Confusinated:

okay thank you